# MotionCtrl: A Unified and Flexible Motion Controller for Video Generation https://wzhouxiff.github.io/projects/MotionCtrl/

### v1

@@ Abstract
Motions in a video primarily consist of camera motion, induced by camera movement, and object motion, resulting from object movement. Accurate control of both camera and object motion is essential for video generation. However, existing works either mainly focus on one type of motion or do not clearly distinguish between the two, limiting their control capabilities and diversity. Therefore, this paper presents MotionCtrl, a unified and flexible motion controller for video generation designed to effectively and independently control camera and object motion. The architecture and training strategy of MotionCtrl are carefully devised, taking into account the inherent properties of camera motion, object motion, and imperfect training data. Compared to previous methods, MotionCtrl offers three main advantages: 1) It effectively and independently controls camera motion and object motion, enabling more fine-grained motion control and facilitating flexible and diverse combinations of both types of motion. 2) Its motion conditions are determined by camera poses and trajectories, which are appearance-free and minimally impact the appearance or shape of objects in generated videos. 3) It is a relatively generalizable model that can adapt to a wide array of camera poses and trajectories once trained. Extensive qualitative and quantitative experiments have been conducted to demonstrate the superiority of MotionCtrl over existing methods.
@@
Figure 1: Control Results of MotionCtrl. MotionCtrl is capable of controlling both camera motion and object motion in videos produced by a video generation model. It can also simultaneously control both types of motion within the same video. We highly encourage readers to check our project page for video results, which cannot be well demonstrated by still images.
[…]
1 Introduction
@@
Video generation, such as text-to-video (T2V) generation [10, 23, 32, 9, 4, 5] aims to produce diverse and high-quality videos that conform to given text prompts. Unlike image generation [21, 17, 6, 34, 18, 19], which focuses on generating a single image, video generation necessitates the creation of consistent and fluent motion among a sequence of generated images. Consequently, motion control plays a significantly crucial role in video generation, yet it has received limited attention in recent research.
@@
In a video, there are primarily two types of motion: global motion induced by camera movements and local motion resulting from object movements (examples are referred to the zoom out camera poses and swaying rose in Figure 1 (c)). It should be noted that these two motions will be consistently referred to as camera motion and object motion throughout the paper, respectively. However, most previous works related to motion control in video generation either primarily focus on one of the motions or lack a clear distinction between these two types of motion. For instance, AnimateDiff [8], Gen-2 [7], and PikaLab [2] mainly execute or trigger camera motion control using independent LoRA [11] models or extra camera parameters (such as “-camera zoom in” in PikaLab [2]). VideoComposer [25] and DragNUWA [28] implement both camera motion and object motion using the same conditions: motion vector in VideoComposer [25] and trajectory in DragNUWA [28]. The lack of clear distinction between these two motions prevents these approaches from achieving fine-grained and diverse motion control in video generation.
@@
In this paper, we design a specialized architecture and training strategy for MotionCtrl to address the aforementioned challenges. First, we construct MotionCtrl with two modules: the Camera Motion Control Module (CMCM) and Object Motion Control Module (OMCM), specifically tailored for camera motion and object motion according to their properties. Both CMCM and OMCM cooperate with the existing video generation model. Noted that in this work, we adopt VideoCrafter1 [5], an improved version over LVDM [9] as the video generation model, and refer it as LVDM throughout the paper. CMCM temporally fuses a sequence of camera poses into the LVDM [9] via its temporal transformers, enabling the global motion of the generated video to conform to the given camera poses. OMCM spatially fuses the information on object movement into the convolutional layers in LVDM to indicate the spatial position of the object in each generated frame.
In this study, we utilize VideoCrafter[5], a model structurally akin to LVDM[9]. For clarity, we will refer to the video model used in this work as LVDM.”
Training the MotionCtrl requires video clips with complete annotations of captions, camera poses, and object movement trajectories, which are currently unavailable and quite difficult to construct. Benefiting from the careful-designed architecture that depends on a large-scale pre-trained video diffusion model equipped with two adapter-like [29, 15] CMCM and OMCM modules, we can train the two modules independently and thus compromise the datasets to one video dataset with the annotations of captions and camera poses and another video dataset with the annotations of captions and object movement trajectories. To this end, we first introduce an augmented-Realestate10k dataset, which exploits Blip2 [13] to generate captions for each sample from Realestate10k [33]. This video dataset with both caption and camera pose annotations is adopted to train the CMCM module. After that, we augment the video from WebVid [3] with object movement trajectories synthesized by a motion segmentation algorithm proposed in ParticleSfM [31]. Thus, we can obtain an augmented WebVid dataset to train the OMCM module. Benefiting from the independent adapter-like training strategy and frozen pre-trained LVDM, we can use one of the camera and object motions or combine both motions to control video generation, enabling fine-grained and flexible motion control.
@@
Through these delicate designs, MotionCtrl demonstrates superiority over previous methods in three aspects: 1) It independently controls camera and object motion, enabling fine-grained adjustments and a variety of motion combinations, as shown in Figure 1. 2) It uses camera poses and trajectories for motion conditions, which do not affect the visual appearance, maintaining the objects’ natural look in videos. For instance, our MotionCtrl generates a video with a camera motion that closely reflects the reference video, offering a realistic Eiffel Tower, as seen in Figure 4 (b). In contrast, VideoComposer [25] relies on dense motion vectors and mistakenly captures a door’s shape of the reference video, resulting in an unnatural Eiffel Tower. 3) MotionCtrl can control a variety of camera movements and trajectories, without the need for fine-tuning each individual camera or object motion.
The main contributions of this work can be summarized as follows:
@@
  • We introduce MotionCtrl, a unified and flexible motion controller for video generation, designed to independently and effectively control camera motion and object motion in generated videos, achieving more fine-grained and diverse motion control.
  • We carefully tailor the architecture and training strategy of MotionCtrl according to the inherent properties of camera motion, object motion, and imperfect training data, effectively achieving fine-grained motion control in video generation.
  • We conduct extensive experiments to demonstrate the superiority of MotionCtrl over previous related methods, both qualitatively and quantitatively.
2 Related Works
Video generation, especially text-to-video (T2V) generation strives to create varied and high-fidelity videos that align with specified textual descriptions. With advancements in diffusion models and the availability of robust computational resources, a plethora of diffusion-based video generation models have emerged[10, 23, 32, 9, 4, 25, 8, 5]. Notably, deploying diffusion models in latent space[19, 9, 4] has enhanced the computational efficiency of video generation. Concurrently, there has been a surge in research focused on controlling motion within generated videos. Many existing approaches learn motion by referencing specific or a series of template videos[26, 8, 27, 30]. While effective at motion control, these methods typically require training a new model for each template or set of templates, which can be limiting.
@@ 2 Related Works
Some efforts aim to achieve more generalized motion control. For instance, VideoComposer [25] introduces motion control via extra provided motion vectors, and DragNUWA [28] suggests video generation conditioned on an initial image, provided trajectories, and text prompts. However, the motion control in these methods is relatively broad and fails to fine-grainedly disentangle the camera and object motion within videos.
@@
Figure 2: MotionCtrl Framework. MotionCtrl extends the Denoising U-Net structure of LVDM with a Camera Motion Control Module (CMCM) and an Object Motion Control Module (OMCM). As illustrated in (b), the CMCM integrates camera pose sequences $RT$ with LVDM’s temporal transformers by appending $RT$ to the input of the second self-attention module and applying a tailored and lightweight fully connected layer to extract the camera pose feature for subsequent processing. The OMCM utilizes convolutional layers and downsamplings to derive multi-scale features from $Trajs$, which are spatially incorporated into LVDM’s convolutional layers to direct object motion. Further given a text prompt, LVDM generates videos from noise that correspond to the prompt, with background and object movements reflecting the specified camera poses and trajectories. The resulting video demonstrates the horse moving along its trajectory and meanwhile, the background moves left, consistent with the camera’s rightward motion.
3 Methodology
3.1 Preliminary
Latent Video Diffusion Model (LVDM). The Latent Video Diffusion Model (LVDM) [9] aims to generate high-quality and diverse videos guided by text prompts. It employs a denoising diffusion model (U-Net [20]) in the latent space for space and time efficiency. Consequently, it constructs a lightweight 3D autoencoder, comprising an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, to encode raw videos into the latent space and reconstruct the denoised latent features back into videos, respectively. Its denoising U-Net (denoted as $\mathcal{\epsilon_{\theta}}$) is constructed with a sequence of blocks that consist of convolutional layers, spatial transformers, and temporal transformers (shown in Figure 2). It is optimized using a noise-prediction loss:
$\mathcal{L}=\mathbb{E}_{z_{0},c,\epsilon\sim\mathcal{N}(0,\mathit{I}),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c)\rVert_{2}^{2}\right],$ (1)
@@ 3.1 Preliminary
$z_{t}=\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon,~{}\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{t},$ (2)
@@ 3.2 MotionCtrl
Figure 2 illustrates the framework of MotionCtrl. To achieve disentanglement between camera motion and object motion, and enable independent control of these two types of motion, MotionCtrl comprises two main components: a Camera Motion Control Module (CMCM) and an Object Motion Control Module (OMCM). Taking into account the global property of camera motion and the local property of object motion, CMCM interacts with the temporal transformers in LVDM, while OMCM spatially cooperates with the convolutional layers in LVDM. Furthermore, we employ multiple training steps to adapt MotionCtrl to the absence of training data that contains high-quality video clips accompanied by captions, camera poses, and object movement trajectories. In the following subsections, we will provide a detailed description of CMCM and OMCM along with their corresponding training datasets and training strategies.
3.2.1 Camera Motion Control Module (CMCM)
@@
The CMCM is a lightweight module constructed with several fully connected layers. Since the camera motions are global transformations between frames in a video, CMCM cooperates with LVDM [9] via its temporal transformers. Typically, the temporal transformers in LVDM comprise two self-attention modules and facilitate temporal information fusion between video frames. To minimize the impact on LVDM’s generative performance, CMCM only involves the second self-attention module in the temporal transformers. Specifically, CMCM takes a sequence of camera poses $RT=\{RT_{0},RT_{1},\dots,RT_{L-1}\}$ as input. In this paper, the camera pose is represented by its $3\times 3$ rotation matrix and $3\times 1$ translation matrix. Consequently, $RT\in\mathbb{R}^{L\times 12}$, where $L$ denotes the length of the generated video. As depicted in Figure 2 (b), $RT$ is extended to $H\times W\times L\times 12$ before being concatenated with the output of the first self-attention module in the temporal transformer ($y_{t}\in\mathbb{R}^{H\times W\times L\times C}$) along the last dimension, where $H$ and $W$ represent the latent spatial size of the generated video, and $C$ is the number of channels in $y_{t}$. The concatenated results are then projected back to the size of ${H\times W\times L\times C}$ using a fully connected layer before being fed into the second self-attention module in the temporal transformer.
3.2.2 Object Motion Control Module (OMCM)
@@
As depicted in Figure 2, MotionCtrl controls the object motion of the generated video using trajectories ($Trajs$). Typically, a trajectory can be represented as a sequence of spatial positions $\{(x_{0},y_{0}),(x_{1},y_{1}),\dots,(x_{L-1},y_{L-1})\}$, where $(x_{i},y_{i}),i\in[0,T-1]$ indicates that the trajectory passes through the $i_{th}$ frame at the spatial position $(x,y)$. While implemented, we tend transform the trajectory $\{(x_{0},y_{0}),(x_{1},y_{1}),\dots,(x_{L-1},y_{L-1})\}$ into $\{(0,0),(u_{1},v_{1}),\dots,(u_{L-1},v_{L-1})\}$, where
$u_{i}=x_{i}-x_{i-1},v_{i}=y_{i}-y_{i-1},i>1.$ (3)
@@
Denoted that the other spatial positions that the trajectories do not pass are described as $(0,0)$ and $Trajs\in\mathbb{R}^{L\times H\times W\times 2}$.
@@
The OMCM spatially incorporates $Trajs$ into the LVDM generation process, as these trajectories represent the spatial positions of the object in each frame. The OMCM’s architecture, highlighted in the purple block of Figure 2, consists of multiple convolutional layers combined with downsampling operations. It extracts multi-scale features from the $Trajs$ and corresponding adds them to the input of the LVDM’s convolutional layers. Drawing inspiration from previous works such as ControlNet[29] and T2I-Adapter [15], the trajectories are only applied to the encoder of the Denoising U-Net to balance the generated video quality with the ability to control object motion.
3.2.3 Training Strategy and Data Construction

### v2

@@ Abstract.
Motions in a video primarily consist of camera motion, induced by camera movement, and object motion, resulting from object movement. Accurate control of both camera and object motion is essential for video generation. However, existing works either mainly focus on one type of motion or do not clearly distinguish between the two, limiting their control capabilities and diversity. Therefore, this paper presents MotionCtrl, a unified and flexible motion controller for video generation designed to effectively and independently control camera and object motion. The architecture and training strategy of MotionCtrl are carefully devised, taking into account the inherent properties of camera motion, object motion, and imperfect training data. Compared to previous methods, MotionCtrl offers three main advantages: 1) It effectively and independently controls camera motion and object motion, enabling more fine-grained motion control and facilitating flexible and diverse combinations of both types of motion. 2) Its motion conditions are determined by camera poses and trajectories, which are appearance-free and minimally impact the appearance or shape of objects in generated videos. 3) It is a relatively generalizable model that can adapt to a wide array of camera poses and trajectories once trained. Extensive qualitative and quantitative experiments have been conducted to demonstrate the superiority of MotionCtrl over existing methods. Project page: https://wzhouxiff.github.io/projects/MotionCtrl/.
AIGC, video generation, motion control
journalyear: 2024copyright: acmlicensedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657518isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Computer vision
@@
Figure 1. Control Results of MotionCtrl. MotionCtrl is capable of controlling both camera motion and object motion in videos produced by a video generation model. It can also simultaneously control both types of motion within the same video. We highly encourage readers to check our project page for video results, which cannot be well demonstrated by still images.
1. Introduction
@@
Video generation, such as text-to-video (T2V) generation (Ho et al., 2022; Singer et al., 2022; Zhou et al., 2022; He et al., 2022; Blattmann et al., 2023b; Chen et al., 2023b) aims to produce diverse and high-quality videos that conform to given text prompts. Unlike image generation (Saharia et al., 2022; Ramesh et al., 2021; Ding et al., 2021; Zhou et al., 2021; Ramesh et al., 2022; Rombach et al., 2022), which focuses on generating a single image, video generation necessitates the creation of consistent and fluent motion among a sequence of generated images. Consequently, motion control plays a significantly crucial role in video generation, yet it has received limited attention in recent research.
@@
In a video, there are primarily two types of motion: global motion induced by camera movements and local motion resulting from object movements (examples are referred to the zoom out camera poses and swaying rose in Fig. 1 (c)). It should be noted that these two motions will be consistently referred to as camera motion and object motion throughout the paper, respectively. However, most previous works related to motion control in video generation either primarily focus on one of the motions or lack a clear distinction between these two types of motion. For instance, AnimateDiff (Guo et al., 2023), Gen-2 (Esser et al., 2023), and PikaLab (pik, [n. d.]) mainly execute or trigger camera motion control using independent LoRA (Hu et al., 2021) models or extra camera parameters (such as ”-camera zoom in” in PikaLab (pik, [n. d.])). VideoComposer (Wang et al., 2023) and DragNUWA (Yin et al., 2023a) implement both camera motion and object motion using the same conditions: motion vector in VideoComposer (Wang et al., 2023) and trajectory in DragNUWA (Yin et al., 2023a). The lack of clear distinction between these two motions prevents these approaches from achieving fine-grained and diverse motion control in video generation.
@@
To address the aforementioned challenges, MotionCtrl deploys a delicately designed architecture, training strategy, and curated datasets. MotionCtrl consists of two modules: the Camera Motion Control Module (CMCM) and the Object Motion Control Module (OMCM), each tailored to handle camera motion and object motion characteristics, respectively. Both CMCM and OMCM function as adapter-like modules integrated into existing video generation models. Specifically, CMCM temporally integrates a sequence of camera poses into the video generation model through its temporal transformers, aligning the global motion of the generated video with the provided camera poses. On the other hand, OMCM spatially incorporates information regarding object movement into the convolutional layers of the video generation model, indicating the spatial positioning of objects in each generated frame. Noted that in this study, we utilize VideoCrafter1 (Chen et al., 2023b), an enhanced version of LVDM (He et al., 2022), as the underlying video generation model, which we refer to as LVDM throughout this paper.
Leveraging the delicately designed architecture reliant on a large-scale pre-trained video diffusion model equipped with adapter-like CMCM and OMCM, we can train these modules separately, thereby mitigating the need for a comprehensive dataset containing videos with annotations of captions, camera poses, and object movement trajectories. Consequently, we achieve MotionCtrl with two datasets: one contains annotations of captions and camera poses, and another comprises annotations of captions and object movement trajectories. Specifically, we introduce the augmented-Realestate10k dataset, originally annotated with camera movement information. We further enhance this dataset by generating captions using Blip2 (Li et al., 2023), rendering it suitable for training camera motion control in video generation. Additionally, we augment videos sourced from WebVid (Bain et al., 2021) with object movement trajectories synthesized using the motion segmentation algorithm proposed in ParticleSfM (Zhao et al., 2022). Alongside their original annotated captions, the augmented-WebVid dataset becomes conducive to learning object motion control in video generation. By sequentially and respectively training CMCM and OMCM with these two annotated datasets, our MotionCtrl framework achieves the capability to independently or jointly control camera and object motion within a unified video generation model. This approach enables relatively fine-grained and flexible motion control, empowering users with enhanced control over the generated videos.
@@
Through these delicate designs, MotionCtrl demonstrates superiority over previous methods in three aspects: 1) It independently controls camera and object motion, enabling fine-grained adjustments and a variety of motion combinations, as shown in Fig. 1. 2) It uses camera poses and trajectories for motion conditions, which do not affect the visual appearance, maintaining the objects’ natural look in videos. For instance, our MotionCtrl generates a video with a camera motion that closely reflects the reference video, offering a realistic Eiffel Tower, as seen in Fig. 4 (b). In contrast, VideoComposer (Wang et al., 2023) relies on dense motion vectors and mistakenly captures a door’s shape of the reference video, resulting in an unnatural Eiffel Tower. 3) MotionCtrl can control a variety of camera movements and trajectories, without the need for fine-tuning each individual camera or object motion.
@@
The main contributions of this work can be summarized as follows: (1) We introduce MotionCtrl, a unified and flexible motion controller for video generation, designed to independently or jointly control camera motion and object motion in generated videos, achieving more fine-grained and diverse motion control. (2) We carefully tailor the architecture and training strategy of MotionCtrl according to the inherent properties of camera motion, object motion, and imperfect training data, effectively achieving fine-grained motion control in video generation. (3) We conduct extensive experiments to demonstrate the superiority of MotionCtrl over previous related methods, both qualitatively and quantitatively.
2. Related Works
Early research in video generation primarily relied on Generative Adversarial Networks (GANs) or Variational Autoencoders (VAEs) (Vondrick et al., 2016; Tulyakov et al., 2018; Wang et al., 2019; Skorokhodov et al., 2022; Wang et al., 2019; Saito et al., 2017). However, in recent years, with the remarkable capacity demonstrated by diffusion models (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022) in image generation, video generation research has shifted towards utilizing diffusion models. By further incorporating with text (Ho et al., 2022; Singer et al., 2022; Zhou et al., 2022; He et al., 2022; Blattmann et al., 2023b; Wang et al., 2023; Guo et al., 2023; Chen et al., 2023b) or image (Yin et al., 2023b; Blattmann et al., 2023a) guidance, diffusion model can generate high-fidelity videos with specific contents. Particularly, the deployment of diffusion models in latent space (Rombach et al., 2022; He et al., 2022; Blattmann et al., 2023b) has significantly enhanced the computational efficiency of video generation, leading to a surge in downstream research centered on diffusion models. MotionCtrl, for instance, aims to leverage diffusion models for controlling motion in generated videos.
@@ 2. Related Works
In the areas of motion control of generated videos, many existing approaches learn motion by referencing specific or a series of template videos(Wu et al., 2023b; Guo et al., 2023; Wu et al., 2023a; Zhao et al., 2023). While effective at a specific motion control, these methods typically require training a new model for different templates, which can be limiting. Some efforts aim to achieve more generalized motion control (Chen et al., 2023a; Wang et al., 2023; Yin et al., 2023a). For instance, VideoComposer (Wang et al., 2023) introduces motion control via extra provided motion vectors, and DragNUWA (Yin et al., 2023a) suggests video generation conditioned on an initial image, provided trajectories, and text prompts. However, the motion control in these methods is relatively broad and fails to fine-grainedly disentangle the camera and object motion within videos.
@@
Figure 2. MotionCtrl Framework. MotionCtrl extends the Denoising U-Net structure of LVDM with a Camera Motion Control Module (CMCM) and an Object Motion Control Module (OMCM). As illustrated in (b), the CMCM integrates camera pose sequences $RT$ with LVDM’s temporal transformers by appending $RT$ to the input of the second self-attention module and applying a tailored and lightweight fully connected layer to extract the camera pose feature for subsequent processing. The OMCM utilizes convolutional layers and downsamplings to derive multi-scale features from $Trajs$, which are spatially incorporated into LVDM’s convolutional layers to direct object motion. Further given a text prompt, LVDM generates videos from noise that correspond to the prompt, with background and object movements reflecting the specified camera poses and trajectories. The resulting video demonstrates the horse moving along its trajectory and meanwhile, the background moves left, consistent with the camera’s rightward motion.
3. Methodology
3.1. Preliminary
The Latent Video Diffusion Model (LVDM) (He et al., 2022) aims to generate high-quality and diverse videos guided by text prompts. It employs a denoising diffusion model (U-Net (Ronneberger et al., 2015)) in the latent space for space and time efficiency. Consequently, it constructs a lightweight 3D autoencoder, comprising an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, to encode raw videos into the latent space and reconstruct the denoised latent features back into videos, respectively. Its denoising U-Net (denoted as $\mathcal{\epsilon_{\theta}}$) is constructed with a sequence of blocks that consist of convolutional layers, spatial transformers, and temporal transformers (shown in Fig. 2). It is optimized using a noise-prediction loss:
(1) $\mathcal{L}=\mathbb{E}_{z_{0},c,\epsilon\sim\mathcal{N}(0,\mathit{I}),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c)\rVert_{2}^{2}\right],$
@@ 3.1. Preliminary
(2) $z_{t}=\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon,~{}\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{t},$
@@ 3.2. MotionCtrl
Fig. 2 illustrates the framework of MotionCtrl. To achieve disentanglement between camera motion and object motion, and enable independent control of these two types of motion, MotionCtrl comprises two main components: a Camera Motion Control Module (CMCM) and an Object Motion Control Module (OMCM). Taking into account the global property of camera motion and the local property of object motion, CMCM interacts with the temporal transformers in LVDM, while OMCM spatially cooperates with the convolutional layers in LVDM. Furthermore, we employ multiple training steps to adapt MotionCtrl to the absence of training data that contains high-quality video clips accompanied by captions, camera poses, and object movement trajectories. In the following subsections, we will provide a detailed description of CMCM and OMCM along with their corresponding training datasets and training strategies.
3.2.1. Camera Motion Control Module (CMCM)
@@
The CMCM is a lightweight module constructed with several fully connected layers. Since the camera motions are global transformations between frames in a video, CMCM cooperates with LVDM (He et al., 2022) via its temporal transformers. Typically, the temporal transformers in LVDM comprise two self-attention modules and facilitate temporal information fusion between video frames. To minimize the impact on LVDM’s generative performance, CMCM only involves the second self-attention module in the temporal transformers. Specifically, CMCM takes a sequence of camera poses $RT=\{RT_{0},RT_{1},\dots,RT_{L-1}\}$ as input. In this paper, the camera pose is represented by its $3\times 3$ rotation matrix and $3\times 1$ translation matrix. Consequently, $RT\in\mathbb{R}^{L\times 12}$, where $L$ denotes the length of the generated video. As depicted in Fig. 2 (b), $RT$ is extended to $H\times W\times L\times 12$ before being concatenated with the output of the first self-attention module in the temporal transformer ($y_{t}\in\mathbb{R}^{H\times W\times L\times C}$) along the last dimension, where $H$ and $W$ represent the latent spatial size of the generated video, and $C$ is the number of channels in $y_{t}$. The concatenated results are then projected back to the size of ${H\times W\times L\times C}$ using a fully connected layer before being fed into the second self-attention module in the temporal transformer.
3.2.2. Object Motion Control Module (OMCM)
@@
As depicted in Fig. 2, MotionCtrl controls the object motion of the generated video using trajectories ($Trajs$). Typically, a trajectory is represented as a sequence of spatial positions $\{(x_{0},y_{0}),(x_{1},y_{1}),\dots,(x_{L-1},y_{L-1})\}$, where $(x_{i},y_{i}),i\in[0,L-1]$ indicates that the trajectory passes through the $i_{th}$ frame at the spatial position $(x,y)$. Particularly, $x\in[0,\hat{W})$ and $y\in[0,\hat{H})$, where $\hat{H}$ and $\hat{W}$ are the height and width of $z_{T}$, respectively. To explicitly expose the moving speed of the object, we represent $Trajs$ as
$\{(0,0),(u_{(x_{1},y_{1})},v_{(x_{1},y_{1})}),\dots,(u_{(x_{L-1},y_{L-1})},v_{(x_{L-1},y_{L-1})})\}$, where
(3) $u_{(x_{i},y_{i})}=x_{i}-x_{i-1};v_{(x_{i},y_{i})}=y_{i}-y_{i-1};0<i<L.$
@@
Denoted that the first frame and the other spatial positions in the subsequent frames that the trajectories do not pass are described as $(0,0)$. Finally, $Trajs\in\mathbb{R}^{L\times\hat{H}\times\hat{W}\times 2}$.
@@
$Trajs$ is injected into LVDM with OMCM, which is highlighted in the purple block of Fig. 2. OMCM consists of multiple convolutional layers combined with downsampling operations. It extracts multi-scale features from the $Trajs$ and correspondingly adds them to the input of the LVDM’s convolutional layers. Drawing inspiration from T2I-Adapter (Mou et al., 2024), the trajectories are only applied to the encoder of the Denoising U-Net to balance the quanlity of the generated video with the ability of object motion control.
3.2.3. Training Strategy and Data Construction